# Sylvester-Preconditioned Adaptive-Rank Implicit Time Integrators for Advection-Diffusion Equations with Variable Coefficients

### v2

@@ C.3 DIRK Generalization to Multi-Rank Advection-Diffusion Coefficients
$\mathbf{F}^{(k)}-a_{kk}\Delta t\left(\underbrace{\sum_{i=1}^{\ell_{x}}T_{1,i}\mathbf{F}^{(k)}{\Phi_{i}^{2,x}}^{\top}+\sum_{j=1}^{\ell_{y}}\Phi_{j}^{1,y}\mathbf{F}^{(k)}{T_{2,j}}^{\top}+\sum_{k=1}^{k_{x}}T_{3,k}\mathbf{F}^{(k)}{\Sigma_{k}^{2,x}}^{\top}+\sum_{l=1}^{k_{y}}\Sigma_{l}^{1,y}\mathbf{F}^{(k)}{T_{4,l}}^{\top}}_{\mathscr{L}(\mathbf{F}^{(k)})}\right)=\mathbf{B}^{(k)},\quad k=1,2,\ldots,s$ (56)
@@
Tolerances $\epsilon_{\kappa}$, $\epsilon_{\texttt{GMRES}}$, $\epsilon$, ${\epsilon}_{\text{tol}}$;
@@
1 Compute operators $\{A_{1,i}\}_{i=1}^{\ell_{x}}$, $\{A_{2,j}\}_{j=1}^{\ell_{y}}$, $\{A_{3,k}\}_{k=1}^{k_{x}}$, $\{A_{4,l}\}_{l=1}^{k_{y}}$ according to Eq. (57);
@@ 7        Orthogonalize and truncate to tolerance $\epsilon_{\kappa}$;
12               Compute $\tilde{{\bf B}}^{(k)}=\tilde{{\bf B}}^{(1)}+\Delta t\sum_{\ell=1}^{k-1}\frac{a_{k\ell}}{a_{\ell\ell}}\left({\bf S}^{(\ell)}-\tilde{{\bf B}}^{(\ell)}\right)$;
@@
17        Compute {_$\{\_,R_{U}\}=\texttt{QR}\left([U_{1},U_{1,\epsilon},T_{1,1:\ell_{x}}U_{1,\epsilon},\Phi_{1:\ell_{y}}^{1,y}U_{1,\epsilon},T_{3,1:k_{x}}U_{1,\epsilon},\Sigma_{1:k_{y}}^{1,y}U_{1,\epsilon}]\right)$ and
18        {_$\{\_,R_{V}\}=\texttt{QR}\left([V_{1},V_{1,\epsilon},\Phi_{1:\ell_{x}}^{2,x}V_{1,\epsilon},T_{2,1:\ell_{y}}V_{1,\epsilon},\Sigma_{1:k_{x}}^{2,x}V_{1,\epsilon},T_{4,1:k_{y}}V_{1,\epsilon}]\right)$ ;
@@ Appendix D Three-Dimensional Formulation
We outline our approach to 3D next. In 3D, we replace the SVD algorithm with a tensor Tucker decomposition, specifically tailored for the three (and higher) dimension tensor-product grids.
@@ D.2 GMRES-ACS Solution of the Reduced GSE in 3D
[…]

### v3

@@ C.3 DIRK Generalization to Multi-Rank Advection-Diffusion Coefficients
$\mathbf{F}^{(k)}-a_{kk}\Delta t\left(\underbrace{\sum_{i=1}^{\ell_{x}}T_{1,i}\mathbf{F}^{(k)}{\Phi_{i}^{2,x}}^{\top}+\sum_{j=1}^{\ell_{y}}\Phi_{j}^{1,y}\mathbf{F}^{(k)}{T_{2,j}}^{\top}+\sum_{k=1}^{k_{x}}T_{3,k}\mathbf{F}^{(k)}{\Sigma_{k}^{2,x}}^{\top}+\sum_{l=1}^{k_{y}}\Sigma_{l}^{1,y}\mathbf{F}^{(k)}{T_{4,l}}^{\top}}_{\mathscr{L}(\mathbf{F}^{(k)})}\right)=\mathbf{B}^{(k)},\qquad k=1,2,\ldots,s$ (56)
@@
1
21exCompute operators $\{A_{1,i}\}_{i=1}^{\ell_{x}}$, $\{A_{2,j}\}_{j=1}^{\ell_{y}}$, $\{A_{3,k}\}_{k=1}^{k_{x}}$, $\{A_{4,l}\}_{l=1}^{k_{y}}$ according to Eq. (57);
@@ 8    Orthogonalize and truncate to tolerance $\epsilon_{\kappa}$;
[…]
@@
18    Compute $\{\_,R_{U}\}=\texttt{QR}\left([U_{1},U_{1,\epsilon},T_{1,1:\ell_{x}}U_{1,\epsilon},\Phi_{1:\ell_{y}}^{1,y}U_{1,\epsilon},T_{3,1:k_{x}}U_{1,\epsilon},\Sigma_{1:k_{y}}^{1,y}U_{1,\epsilon}]\right)$ and
19    $\{\_,R_{V}\}=\texttt{QR}\left([V_{1},V_{1,\epsilon},\Phi_{1:\ell_{x}}^{2,x}V_{1,\epsilon},T_{2,1:\ell_{y}}V_{1,\epsilon},\Sigma_{1:k_{x}}^{2,x}V_{1,\epsilon},T_{4,1:k_{y}}V_{1,\epsilon}]\right)$ ;
@@ Appendix D Three-Dimensional Formulation
We outline the approach to 3D next. In 3D, we replace the SVD algorithm with a tensor Tucker decomposition, specifically tailored for the three (and higher) dimension tensor-product grids.
@@ D.2 GMRES-ACS Solution of the Reduced GSE in 3D
$\mathbf{F}_{0}=S_{0}\times_{1}U_{0}\times_{2}V_{0}\times_{3}W_{0}\xrightarrow{\text{evolve in time}}\mathbf{F}_{1}=S_{1}\times_{1}U_{1}\times_{2}V_{1}\times_{3}W_{1},$ (65)
@@
Appendix E Technical Details of the Hypothetical Multigrid Comparison
We provide here a brief explanation of the assumptions involved in the performance comparison between a hypothetical optimally scaling full-rank multigrid method and the adaptive-rank approach. Optimal multigrid methods exhibit a computational scaling of $\mathcal{O}(N^{d}\log N)$ [21]. Let us denote the wall-clock time (WCT) of a multigrid method as:
$WCT_{\mathrm{MG}}=C_{\mathrm{MG}}N^{d}\log N.$
From Fig. 1-a, the wall-clock time of the adaptive-rank method may be written as:
$WCT_{\mathrm{LR}}=C_{\mathrm{LR}}(0.0023\,N+55.2734).$
The speedup of the adaptive-rank method vs. multigrid is given by the ratio:
$\frac{WCT_{\mathrm{MG}}}{WCT_{\mathrm{LR}}}=C_{\mathrm{WCT}}\frac{N^{3}\log N}{0.0023\,N+55.2734},$ (68)
where $C_{\mathrm{WCT}}=1.2\times 10^{-5}$ is found assuming that both methods achieve comparable wall-clock times for $N=100$. This formula is used to populate Table 1.